### v1

@@ 5.2 Main Auditing Results
PANORAMIA $2.508$ $2.83$ $0.32$
ResNet101_E20 PANORAMIARM;RN 00 $0.42$ $0.42$
O (1) RM;RN - - - $0.52$
PANORAMIA $2.508$ $3.15$ $0.64$
ResNet101_E50 PANORAMIARM;RN 00 $0.61$ $0.61$
O (1) RM;RN - - - $0.81$
PANORAMIA $2.508$ $3.47$ $0.962$
ResNet101_E100 PANORAMIARM;RN 00 $1.03$ $1.03$
O (1) RM;RN - - - $1.40$
PANORAMIA $2.01$ $2.50$ $0.49$
CNN_E50 PANORAMIARM;RN 00 - 0.76
O (1) RM;RN - - - 0.99
PANORAMIA $2.01$ $3.03$ $1.02$
CNN_E100 PANORAMIARM;RN 00 - 1.26
O (1) RM;RN - - - 1.53
PANORAMIA $3.78$ $3.47$ 00 -
GPT2_E12 PANORAMIARM;RN 00 $0.30$ $0.30$
O (1) RM;RN - - - $1.54$
PANORAMIA $3.78$ $4.07$ $0.29$
GPT2_E46 PANORAMIARM;RN 00 $2.37$ $2.37$
O (1) RM;RN - - - $4.12$
PANORAMIA $3.78$ $5.06$ $1.28$
GPT2_E92 PANORAMIARM;RN 00 $3.45$ $3.45$
O (1) RM;RN - - - $5.43$
PANORAMIA $2.37$ $2.47$ $0.10$
MLP_E10 O (1) RM;RN - - - $0$
PANORAMIA $2.37$ $2.71$ $0.34$
MLP_E100 O (1) RM;RN - - - $0.23$
PANORAMIA $1.25$ $1.62$ $0.37$
MLP_E100_half PANORAMIARM;RN 00 $0.64$ $0.64$
O (1) RM;RN - - - $0.22$
Table 4: Privacy audits on different target models.
LABEL:fig:eval:comparison_precision shows the precision of $b$ and PANORAMIA at different levels of recall, and LABEL:fig:eval:comparison the corresponding value of $\{c+\epsilon\}_{lb}$ (or $c_{\textrm{lb}}$ for $b$). Dashed lines show the maximum value of $\{c+\epsilon\}_{lb}$/$c_{\textrm{lb}}$ achieved (Fig. LABEL:fig:eval:comparison) (returned by PANORAMIA), and the precision implying these values at different recalls (Fig. LABEL:fig:eval:comparison_precision). Table 4 summarizes those $\{c+\epsilon\}_{lb}$/$c_{\textrm{lb}}$ values, as well as the $\epsilon$ measured by existing approaches. We make two key observations.
@@
First, the best prior method (whether RM;RN or O(1)) measures a larger privacy loss ($\tilde{\epsilon}\leq\epsilon$), except on tabular data. Those surprising results are likely due to PANORAMIA’s use of both raw data and the target model’s loss in an ML MIA model, whereas O(1) uses a threshold value on the loss only. Overall, these results empirically confirm the strength of $b$, as we do not seem to spuriously assign differences between ${\mathcal{G}}$ and ${\mathcal{D}}$ to our privacy loss proxy $\tilde{\epsilon}$. We also note that O(1) tends to perform better, due to its ability to rely on non-member detection, which improves the power of the statistical test at equal data sizes. Such tests are not available in PANORAMIA given our one-sided closeness definition for ${\mathcal{G}}$ (see §4), and we keep that same one-sided design for RM;RN for comparison’s sake.
@@
Second, the values of $\tilde{\epsilon}$ measured by PANORAMIA are close to those of the methods against which we compared. In particular, despite a more restrictive adversary model (i.e., no non-member data, no control over the training process, and no shadow model training), PANORAMIA is able to detect meaningful amounts of privacy loss, comparable to that of state-of-the-art methods! For instance, on a non-overfitted CIFAR-10 model (E20), PANORAMIA detects a privacy loss of $0.32$, while using real non-member (RM;RN) data yields $0.42$, and controlling the training process $O(1)$ gets $0.52$. The relative gap gets even closer on models that reveal more about their training data. Indeed, for the most over-fitted model (E100), $\tilde{\epsilon}=0.96$ is very close to RM;RN ($\epsilon=1.0$) and O(1) ($\epsilon=1.4$). This also confirms that the leakage detected by PANORAMIA on increasingly over-fitted models does augment, which is confirmed by prior state-of-the-art methods. For instance, NLP models’ $\tilde{\epsilon}$ goes from $0.18$ to $1.28$ ($1.54$ to $5.43$ for O(1)), and tabular data MLPs from $0.1$ to $0.34$ (00 to $0.23$ for O(1)).
5.3 Detecting Controlled Variations in Privacy Loss
Models of varying complexity: Carlini et al. (2021) have shown that larger models tend to have bigger privacy losses. To confirm this, we conducted an audit of ML models with varying numbers of parameters, from a $\approx 4M$ parameters Wide ResNet-28-2, to a $25.5M$ parameters ResNet-50, and a $44.5M$ parameters ResNet-101. LABEL:subfig:var_complexities shows that PANORAMIA does detect increasing privacy leakage, with $\tilde{\epsilon}_{wide-resnet}\leq\tilde{\epsilon}_{resnet50}\leq\tilde{\epsilon}_{resnet101}$.
ResNet18 $\epsilon=\infty$ PANORAMIA RM;GN 2.508 3.6698 1.161 -
O (1) RM;RN - - - 1.565
ResNet18 $\epsilon=15$ PANORAMIA RM;GN 2.508 3.5707 1.062 -
O (1) RM;RN - - - 1.22
ResNet18 $\epsilon=10$ PANORAMIA RM;GN 2.508 2.8 0.3 -
O (1) RM;RN - - - 0.14
ResNet18 $\epsilon=2$ PANORAMIA RM;GN 2.508 2.065 0 -
O (1) RM;RN - - - 0.08
ResNet18 $\epsilon=1$ PANORAMIA RM;GN 2.508 1.982 0 -
O (1) RM;RN - - - 0
Table 5: $\epsilon$-DP ResNet18 on CIFAR-10.
@@ 5.3 Detecting Controlled Variations in Privacy Loss
DP models: Another avenue to varying privacy leakage is by training DP models with diverse values of $\epsilon$. We evaluate PANORAMIA on DP ResNet-18 models on CIFAR10, with $\epsilon$ values shown on Table 5 and LABEL:subfig:resnet18-dp-fig. The hyper-parameters were tuned independently for the highest train accuracy (see Appendix C.6 for more results and discussion). As the results show, neither PANORAMIA or O(1) detect privacy loss on the most private models ($\epsilon=1,2$). At higher values of $\epsilon=10,15$ (i.e., less private models) and $\epsilon=\infty$ (i.e., non-private model) PANORAMIA does detect an increasing level of privacy leakage with $\tilde{\epsilon}_{\epsilon=10}<\tilde{\epsilon}_{\epsilon=15}<\tilde{\epsilon}_{\epsilon=\infty}$. In this regime, the O(1) approach detects a larger (except for $\epsilon=10$, in which PANORAMIA surprisingly detects a slightly larger value), though comparable, amount of privacy loss.
Appendix A Proofs
@@ A.1 Proof of Proposition 1
${\mathbb{P}}_{S,X,T^{b}}\Big{[}\sum_{i=1}^{m}T^{b}_{i}\cdot S_{i}\geq v\ |\ T^{b}=t^{b}\Big{]}$
[…]
@@ A.2 Proof of Proposition 2
${\mathbb{P}}_{S,X,T^{a}}\Big{[}\sum_{i=1}^{m}T^{a}_{i}\cdot S_{i}\geq v\ |\ T^{a}=t^{a}\Big{]}$
[…]
@@
${\mathbb{P}}\big{[}T^{a}=t^{a}\ |\ S_{<i}=s_{<i},X_{\leq i}=x_{\leq i}\big{]}=$
$={\mathbb{P}}\big{[}T^{a}=t^{a}\ |\ S_{i}=1,S_{<i}=s_{<i},X_{\leq i}=x_{\leq i}\big{]}$
$\ \ \cdot{\mathbb{P}}\big{[}S_{i}=1\ |\ S_{<i}=s_{<i},X_{\leq i}=x_{\leq i}\big{]}$
$+{\mathbb{P}}\big{[}T^{a}=t^{a}\ |\ S_{i}=0,S_{<i}=s_{<i},X_{\leq i}=x_{\leq i}\big{]}$
$\ \ \cdot{\mathbb{P}}\big{[}S_{i}=0\ |\ S_{<i}=s_{<i},X_{\leq i}=x_{\leq i}\big{]},$
@@
Appendix B Experimental Details
@@
B.1 Image data
Target models. We audit target models with the following architectures: a Multi-Label Convolutional Neural Network (CNN) with four layers O’Shea & Nash (2015), and the ResNet101 He et al. (2015). We also include in our analysis, differentially-private models for ResNet18 He et al. (2015) and WideResNet-16-4 Zagoruyko & Komodakis (2017) models as targets, with $\epsilon=1,2,4,6,10,15,20$. The ResNet-based models are trained on CIFAR10 using $50k$ images Krizhevsky et al. (2009) of $32$x32323232 resolution. For all CIFAR10 based classification models (apart from the DP ones), we use a training batch size of 64. The associated test accuracies and epochs are mentioned in Table 2. The Multi-Label CNN is trained on $200k$ images of CelebA Liu et al. (2015) of $128$x128128128128 resolution, training batch-size 32, to predict $40$ attributes associated with each image.

### v2

@@ 5.2 Main Privacy Measurement Results
PANORAMIA $2.44\pm 0.190$ $2.76\pm 0.250$ $0.33\pm 0.21$
ResNet101_E20 PANORAMIA RM;RN $0.00$ $0.421\pm 0.19$ $0.421\pm 0.19$
O (1) RM;RN - - - $0.450\pm 0.150$
PANORAMIA $2.44\pm 0.190$ $3.11\pm 0.18$ $0.67\pm 0.17$
ResNet101_E50 PANORAMIA RM;RN $0.00$ $0.71\pm 0.21$ $0.71\pm 0.21$
O (1) RM;RN - - - $0.79\pm 0.23$
PANORAMIA $2.44\pm 0.190$ $3.60\pm 0.21$ $1.16\pm 0.19$
ResNet101_E100 PANORAMIA RM;RN $0.00$ $1.22\pm 0.190$ $1.22\pm 0.19$
O (1) RM;RN - - - $1.41\pm 0.180$
PANORAMIA $2.44\pm 0.190$ $5.01\pm 0.25$ $2.57\pm 0.23$
WRN-28-10_E150 PANORAMIA RM;RN $0.00$ $2.87\pm 0.20$ $2.87\pm 0.20$
O (1) RM;RN - - - $2.96\pm 0.19$
PANORAMIA $2.44\pm 0.190$ $5.98\pm 0.23$ $3.54\pm 0.23$
WRN-28-10_E300 PANORAMIA RM;RN $0.00$ $3.65\pm 0.22$ $3.65\pm 0.22$
O (1) RM;RN - - - $3.73\pm 0.19$
PANORAMIA $2.44\pm 0.190$ $5.03\pm 0.24$ $2.59\pm 0.22$
ViT_E35 PANORAMIA RM;RN $0.00$ $2.77\pm 0.21$ $2.77\pm 0.21$
O (1) RM;RN - - - $2.86\pm 0.19$
PANORAMIA $2.05\pm 0.21$ $2.81\pm 0.24$ $0.76\pm 0.23$
CNN_E50 PANORAMIA RM;RN $0.00$ $0.91\pm 0.22$ $0.91\pm 0.22$
O (1) RM;RN - - - $1.01\pm 0.18$
PANORAMIA $2.05\pm 0.21$ $3.38\pm 0.24$ $1.33\pm 0.22$
CNN_E100 PANORAMIA RM;RN $0.00$ $1.38\pm 0.21$ $1.38\pm 0.21$
O (1) RM;RN - - - $1.46\pm 0.16$
PANORAMIA $3.62\pm 0.32$ $3.85\pm 0.46$ $0.22\pm 0.37$
GPT2_E37 PANORAMIA RM;RN $0.00$ $1.04\pm 0.42$ $1.04\pm 0.42$
O (1) RM;RN - - - $2.82\pm 0.31$
PANORAMIA $3.62\pm 0.32$ $5.20\pm 0.34$ $1.57\pm 0.45$
GPT2_E75 PANORAMIA RM;RN $0.00$ $2.52\pm 0.36$ $2.52\pm 0.36$
O (1) RM;RN - - - $4.71\pm 0.34$
PANORAMIA $3.62\pm 0.32$ $6.02\pm 0.34$ $2.40\pm 0.53$
GPT2_E150 PANORAMIA RM;RN $0.00$ $3.60\pm 0.41$ $3.60\pm 0.41$
O (1) RM;RN - - - $5.73\pm 0.08$
ResNet18 $\epsilon=\infty$ PANORAMIA $2.44\pm 0.190$ $3.87\pm 0.20$ $1.43\pm 0.21$
O (1) RM;RN - - - $1.471\pm 0.13$
ResNet18 $\epsilon=15$ PANORAMIA $2.44\pm 0.190$ $3.57\pm 0.19$ $1.13\pm 0.19$
O (1) RM;RN - - - $1.20\pm 0.18$
ResNet18 $\epsilon=10$ PANORAMIA $2.44\pm 0.190$ $2.70\pm 0.25$ $0.26\pm 0.22$
O (1) RM;RN - - - $0.28\pm 0.14$
ResNet18 $\epsilon=2$ PANORAMIA $2.44\pm 0.190$ $1.709\pm 0.23$ $0.00$
O (1) RM;RN - - - $0.05\pm 0.12$
ResNet18 $\epsilon=1$ PANORAMIA $2.44\pm 0.190$ $1.412\pm 0.12$ $0.00$
O (1) RM;RN - - - $0.00$
Table 3: Privacy measurements on different target models ($m=10k$). O(1) RM;RN with our ML-based attack for $\epsilon$-DP models shown in LABEL:subfig:panoramia_vs_O1_m_dp_increase.
@@
First, the best prior method (whether RM;RN or O(1)) measures a larger privacy loss ($\tilde{\epsilon}\leq\epsilon$). Overall, these results empirically confirm the strength of $b$, as we do not seem to spuriously assign differences between ${\mathcal{G}}$ and ${\mathcal{D}}$ to our privacy loss proxy $\tilde{\epsilon}$. We also note that O(1) tends to perform better, due to its ability to rely on non-member detection, which improves the power of the statistical test at equal data sizes. Such tests are not available in PANORAMIA given our one-sided closeness definition for ${\mathcal{G}}$ (see §4), and we keep that same one-sided design for RM;RN for the sake of comparison.
@@
Second, the values of $\tilde{\epsilon}$ measured by PANORAMIA are close to those of the methods against which we compared. In particular, despite a more restrictive adversary model (i.e., no non-member data, no control over the training process, and no shadow model training), PANORAMIA is able to detect meaningful amounts of privacy loss, comparable to that of state-of-the-art methods.
For instance, on a non-overfitted CIFAR-10 ResNet101 model (E20), PANORAMIA detects a privacy loss of $0.33$, while using real non-member (RM;RN) data yields $0.42$, and controlling the training process $O(1)$ gets $0.45$. The relative gap gets even closer on models that reveal more about their training data. Indeed, for the most over-fitted model (E100), $\tilde{\epsilon}=1.16$ is very close to RM;RN ($\epsilon=1.22$) and O(1) ($\epsilon=1.41$). This also confirms that the leakage detected by PANORAMIA on increasingly overfitted models does augment, which is confirmed by prior state-of-the-art methods. For instance, NLP models’ $\tilde{\epsilon}$ goes from $0.22$ to $2.40$ ($2.82$ to $5.73$ for O(1)). Section D.6 details results on tabular data, in which PANORAMIA is not able to detect significant privacy loss.
Privacy Measurement of $\epsilon$-DP Models:
Figure 5: ResNet18, CIFAR-10, $\epsilon$-DP for various $\epsilon$ values.
@@
Another avenue to varying privacy leakage is by training Differentially-Private (DP) models with different values of $\epsilon$. We evaluate PANORAMIA on DP ResNet-18 models on CIFAR10, with $\epsilon$ values shown on Table 3 and Figure 5. The hyper-parameters were tuned independently for the highest train accuracy (see Appendix D.9 for more results and discussion). As the results show, neither PANORAMIA nor O(1) detect privacy loss on the most private models ($\epsilon=1,2$). At higher values of $\epsilon=10,15$ (i.e., less private models) and $\epsilon=\infty$ (i.e., non-private model) PANORAMIA does detect an increasing level of privacy leakage with $\tilde{\epsilon}_{\epsilon=10}<\tilde{\epsilon}_{\epsilon=15}<\tilde{\epsilon}_{\epsilon=\infty}$. In this regime, the O(1) approach detects a larger, though comparable, amount of privacy loss.
We also evaluate PANORAMIA on DP (fine-tuned) large language models (see Appendix D.9 and Table 9 for details). We fine-tune GPT2-Small target models with DP-SGD on the WikiText dataset, with various values of $\epsilon$. Neither PANORAMIA nor O(1) are able to measure a positive privacy loss.
5.3 Leveraging More Data to Improve Privacy Measurements
PANORAMIA can leverage much more data for its measurement (up to the whole training set size for training and testing the MIA), while in our setting O(1) is limited by the size of the test set (for non-members). As we have seen in LABEL:subfig:main-test-inc-resnet100, PANORAMIA can leverage larger test set size to measure higher privacy loss values than O(1), up to $2.62$ on the ResNet101 trained for 100 epochs, versus $2.23$ for O(1). It is also important to note that at small amounts of data O(1) measures a larger privacy loss ($\tilde{\epsilon}\leq\epsilon$). These findings provide additional empirical evidence for the robustness of $b$, indicating that we are not erroneously attributing differences between ${\mathcal{G}}$ and ${\mathcal{D}}$ to our privacy loss proxy $\tilde{\epsilon}$. However, in the case of text data (on the GPT-2 trained for 150 epochs), LABEL:subfig:main-test-inc-gpt2, we hit an upper-bound on the power of the hypothesis test which estimates $\{c\!+\!\epsilon\}_{\textrm{lb}}$ using $20k$ test samples. The maximum measurement for $\tilde{\epsilon}$ we can achieve is around $2.59$ with a $20k$ test set, as opposed to O(1), which reaches $5.73$ with $10k$ test set size.
Increasing Test Set Size for DP models: With an increase in the test set size $m$, we demonstrate that we can achieve tighter bounds on the privacy measurement $\tilde{\epsilon}$. LABEL:subfig:panoramia_vs_O1_m_dp_increase shows that as $m$ increases, PANORAMIA (solid lines) can measure higher privacy leakage for various $\epsilon$-DP ResNet18 models. The reason for this is that PANORAMIA can leverage a much higher test set size (up to the whole training set size for training and testing the MIA), in comparison to O(1) in our setting (due to being limited by the number of real non-member data points available).
6 Impact, Limitations and Future Directions
We have introduced a new approach to quantify the privacy leakage from ML models, in settings in which individual data contributors (such as a hospital in a cross-site FL setting or a user of a text auto-complete service) measure the leakage of their own, known partial training data in the final trained model. Our approach does not introduce new attack capabilities (it would likely benefit from future progress on MIA performance though), but can help model providers and data owners assess the privacy leakage incurred by their data, due to inclusion in ML models training datasets. Consequently, we believe that any impact from our proposed work is likely to be positive.
However, PANORAMIA suffers from a major limitation: its privacy measurement is not a lower-bound on the privacy loss (see details in Section 4.2). Providing a proper lower-bound $\epsilon_{lb}$ on privacy leakage in this setting is an important avenue for future work. A promising direction is to devise a way to measure or enforce an upper-bound on $c$, thereby yielding $\epsilon_{lb}$ via $\{c+\epsilon\}_{lb}-c_{ub}$. Despite this shortcoming, we believe that the new measurement setting we introduce in this work is important. Indeed, recent work has shown that MIA evaluation on foundation models suffers from a similar lack of non-member data (since all in distribution data is included in the model) Das et al. (2024); Duan et al. (2024); Meeus et al. (2024). The theory we develop in this paper provides a meaningful step towards addressing privacy measurements in this setting, and provides a more rigorous approach to privacy benchmarks for such models. We also demonstrate that PANORAMIA’s privacy measurements can also be empirically valuable for instance for providing improved measurements with more data (LABEL:fig:eval:test-size-growth).
Acknowledgments
Mathias Lécuyer is grateful for the support of the Natural Sciences and Engineering Research Council of Canada (NSERC) [reference number RGPIN-2022-04469], as well as a Google Research Scholar award. Sébastien Gambs is supported by the Canada Research Chair program, a Discovery Grant from the Natural Sciences and Engineering Research Council of Canada (NSERC) as well as the DEEL Project CRDPJ 537462-18 funded by the NSERC and the Consortium for Research and Innovation in Aerospace in Québec (CRIAQ). This research was enabled by computational support provided by the Digital Research Alliance of Canada (alliancecan.ca). Hadrien Lautraite is supported by National Bank of Canada and a bursary from the Canada Research Chair in Privacy-preserving and Ethical Analysis of Big Data..
Appendix A Notations
Table 4 summarizes the main notations used in the paper.
Notation Description
$f$ the target model to be audited.
${\mathcal{G}}$ the generative model
${\mathcal{D}}$ distribution over auditor samples
$D_{f}$ (subset of the) training set of the target model $f$ from ${\mathcal{D}}$
$m$ $|D_{\textrm{in}}|=|D_{\textrm{out}}|\triangleq m$
$b$ baseline classifier for $D_{\textrm{in}}$ vs. $D_{\textrm{out}}$
Table 4: Summary of notations
{hideIfProofs}
Appendix B Proofs
@@ B.1 Proof of Proposition 1
${\mathbb{P}}_{S,X,T^{b}}\Big{[}\sum_{i=1}^{m}T^{b}_{i}\cdot S_{i}\geq v\ |\ T^{b}=t^{b}\Big{]}\leq\underset{{S^{\prime}\sim\textrm{Bernoulli}(\frac{e^{c}}{1+e^{c}})^{m}}}{{\mathbb{P}}}\Big{[}\sum_{i=1}^{m}t^{b}_{i}\cdot S^{\prime}_{i}\geq v\Big{]}\triangleq\beta^{b}(m,c,v,t^{b})$
@@ B.2 Proof of Proposition 2
${\mathbb{P}}_{S,X,T^{a}}\Big{[}\sum_{i=1}^{m}T^{a}_{i}\cdot S_{i}\geq v\ |\ T^{a}=t^{a}\Big{]}\leq\underset{{S^{\prime}\sim\textrm{Bernoulli}(\frac{e^{c+\epsilon}}{1+e^{c+\epsilon}})^{m}}}{{\mathbb{P}}}\Big{[}\sum_{i=1}^{m}t^{a}_{i}\cdot S^{\prime}_{i}\geq v\Big{]}\triangleq\beta^{a}(m,c,\epsilon,v,t^{a})$
@@
${\mathbb{P}}\big{[}T^{a}=t^{a}\ |\ S_{<i}=s_{<i},X_{\leq i}=x_{\leq i}\big{]}$
$={\mathbb{P}}\big{[}T^{a}=t^{a}\ |\ S_{i}=1,S_{<i}=s_{<i},X_{\leq i}=x_{\leq i}\big{]}\cdot{\mathbb{P}}\big{[}S_{i}=1\ |\ S_{<i}=s_{<i},X_{\leq i}=x_{\leq i}\big{]}$
$+{\mathbb{P}}\big{[}T^{a}=t^{a}\ |\ S_{i}=0,S_{<i}=s_{<i},X_{\leq i}=x_{\leq i}\big{]}\cdot{\mathbb{P}}\big{[}S_{i}=0\ |\ S_{<i}=s_{<i},X_{\leq i}=x_{\leq i}\big{]},$
@@
$\frac{{\mathbb{P}}\big{[}T^{a}=t^{a}\ |\ S_{i}=0,S_{<i}=s_{<i},X_{\leq i}=x_{\leq i}\big{]}}{{\mathbb{P}}\big{[}T^{a}=t^{a}\ |\ S_{i}=1,S_{<i}=s_{<i},X_{\leq i}=x_{\leq i}\big{]}}\geq e^{-\epsilon}.$ (6)
This inequality might seem surprising at first, as we know from the privacy game (Definition 2) that $S$ is independent of $f$, that is $S\perp\!\!\!\!\perp f$. However, these two quantities are not conditionally independent when conditioning on $X$ (where $X$ is the test set for the MIA/baseline), (i.e., $S\not\perp\!\!\!\!\perp f|X$). This is because $X_{i}$ is either a member or non-member based on $S_{i}$. So when $S_{i}=1$, then $f$ was trained on $X_{i}$, which makes $f$ and $S_{i}$ dependent if membership can be detected through $f$. If $f$ is $\epsilon$-DP, we can bound this dependency using DP properties, which is what we do in Equation 6.
@@
$=\frac{{\mathbb{P}}\big{[}X_{i}=x_{i}\ |\ S_{i}=0,S_{<i}=s_{<i},X_{<i}=x_{<i}\big{]}}{{\mathbb{P}}\big{[}X_{i}=x_{i}\ |\ S_{i}=1,S_{<i}=s_{<i},X_{<i}=x_{<i}\big{]}}\cdot\frac{{\mathbb{P}}\big{[}S_{i}=0\ |\ S_{<i}=s_{<i},X_{<i}=x_{<i}\big{]}}{{\mathbb{P}}\big{[}S_{i}=1\ |\ S_{<i}=s_{<i},X_{<i}=x_{<i}\big{]}}$
@@
Appendix C Experimental Details
@@
C.1 Image data
Target models. We audit target models with the following architectures: a Multi-Label Convolutional Neural Network (CNN) with four layers O’Shea & Nash (2015), and the ResNet101 He et al. (2015) as well as a Wide ResNet-28-10  Zagoruyko & Komodakis (2017b), and a Vision Transformer (ViT)-small  Dosovitskiy et al. (2021). We also include in our analysis, differentially-private models for ResNet18 He et al. (2015) and WideResNet-16-4 Zagoruyko & Komodakis (2017a) models as targets, with $\epsilon=1,2,4,6,10,15,20$. The ResNet-based models are trained on CIFAR10 using $50k$ images Krizhevsky (2009) of $32$x32323232 resolution. For ResNet101 CIFAR10-based classification models, we use a training batch size of 32; for WideResNet-28-10, we use batch size 128. The details for training DP models are given in D.9. The associated test accuracies and epochs are mentioned in Table 1. The Multi-Label CNN is trained on $200k$ images of CelebA Liu et al. (2015) of $128$x128128128128 resolution, training batch-size 32, to predict $40$ attributes associated with each image.